Dublin University Rifle Club

Version Date : April 23, 2002

DURC Competition Attendance Report

Fassaroe Air Rifle Open (Stage one)

Distribution

Dr. Trevor West DUCAC Chairman
Aubrey Storey DUCAC Secretary
Terry McAuley DURC President and Director of Sport
Michelle Bennett Physical Recreation Officer
Tony O’Donnell Student Record
Susan Thompson Trinity News

Summary

Venue UCD Rifle Club
Date 10-11 November 2001
Discipline 10m Air Rifle

DURC Attendees

Steven Watterson  Mark Dennehy
Ben Jones  John Keeney
Peter Byrne  Adrian Costigan

Coleman Marcus-Quinn

Scores

Steven Watterson 565
Mark Dennehy 553, 552 (reentry)
Ben Jones 486
John Keeney 438
Peter Byrne 476, 465 (reentry)
Adrian Costigan 457, 427 (reentry)
Coleman Marcus-Quinn 441

Prizes Won

Mark Dennehy Class B, 1st place
Peter Byrne Class D, 3rd place
Coleman Marcus-Quinn Best DURC Novice shooter

Report

DURC sent a small team to compete in the Fassaroe Club’s Air Rifle Open (aka the WAIDAM open) this year, with a total of 7 competitors.

The standard of competition was high, the winning score being 578 from Nicola Murphy of Wilkinstown (an ex-UCD shooter), who recently won silver in the English championship and who has competed internationally representing Ireland many times. Steven shot lower than his previous average, with his 565 beaten into 4th place by one point by Declan Kelly of the Fassaroe Club (566). (Second place was taken by Olympian Rhona Barry with 570).

Unfortunately, most of the DURC shooters were only just recovering from their annual bout of ’flu and so scores were generally down. Mark shot a personal best on Saturday with 553 and then reentered on Sunday hoping to push that further, but instead dropped one point to 552. Peter Byrne shot lower than he had hoped, but still managed to take third place in class D. Novice Colman Marcus-Quinn managed to increase his Personal Best score by 82 points to 441 and took the prize for best DURC novice.

The next competition is our own DURC 10m Air Rifle Open, to be held in UCD on the 24^{th} and 25^{th} of November. We’re hoping to get about 40 competitiors including our own members.

John Keeney and Steven Watterson shooting; Adrian Costigan shooting

Peter Byrne shooting; John Woodward (Wilkinstown) shooting - note the size and distance of the target!

Misé le meas,

Mark Dennehy, DURC Secretary